OpenStudy (anonymous):

What is the average of the 1st 100 counting numbers?

OpenStudy (anonymous):

average =50*101/100 =50.5

OpenStudy (amistre64):

depends on where youre counting from id say

OpenStudy (anonymous):

Why do you multiply 50 to it?

OpenStudy (amistre64):

100/2 = 50

OpenStudy (anonymous):

Please explain further.